the permiter of a rectangle is 164m and the length is 7m more than twice the width, could you show me how to find the width and the length?
Answer by checkley77(12844)

P=2L+2W
L=2W+7
164=2(2W+7)+2W
164=4W+14+2W
6W=164-14
6W=150
W=150/6
W=25 ANS. FOR THE WIDTH.
L=2*25+7
L=50+7=57 ANS. FOR THE LENGTH.
PROOF:
164=2*57=2*25
164=114+50
164=164